OpenStudy (daenio):

Two objects are experiencing a force of gravitational attraction. If you triple the mass of one of the objects and double the distance between their centres, the new force of gravity compared to the old (Fg) will be: Question 3 options: A) 3 Fg B) 1.5 Fg C) 0.75 Fg D) the same

OpenStudy (daenio):

@Shane_B @CarlosGP

OpenStudy (goformit100):

B) 1.5 Fg

OpenStudy (daenio):

That's what I thought too. How did you get B)?

OpenStudy (shane_b):

Check your last post...C was not correct :) I'll look at this one in just a moment

OpenStudy (shane_b):

For this one, you can just write out both scenarios and see what the difference would be:\[F_{g1}=G\frac{m_1m_2}{r^2}\]\[F_{g1}=G\frac{3m_1m_2}{(2r)^2}\]Just looking at what's different you should see that:\[\frac{3}{2^2}=\frac{3}{4}=0.75\]

OpenStudy (daenio):

Hey, so for the last post, is C) correct?

OpenStudy (shane_b):

Yep

OpenStudy (daenio):

Cool, thank you!

OpenStudy (shane_b):

No worries, good luck!

OpenStudy (daenio):

Thanks, hey I have about 2 more questions.

OpenStudy (daenio):

I guess I'll post them in a new thread.

OpenStudy (shane_b):

ok